The stockholders’ equity of TVX Company at the beginning of the day on February 5 follows: Common stock—$25 par value, 150,000 shares authorized, 59,000 shares issued and outstanding $ 1,475,000 Paid-in capital in excess of par value, common stock 525,000 Retained earnings 675,000 Total stockholders’ equity $ 2,675,000 On February 5, the directors declare a 16% stock dividend distributable on February 28 to the February 15 stockholders of record. The stock’s market value is $35 per share on February 5 before the stock dividend. The stock’s market value is $30 per share on February 28. 2. One stockholder owned 450 shares on February 5 before the dividend. Compute the book value per share and total book value of this stockholder’s shares immediately before and after the stock dividend of February 5. (Round your "Book value per share" answers to 3 decimal places.)
Answer #1

Before Stock Dividend:

Book value per share immediately before the stock dividend = $ 2,675,000 / 59,000 shares = $ 45.339

Total book value for the stockholder holding 450 shares = 450 x $ 45.339 = $ 20,402.55

After Stock Dividend :

Total book value remains unchanged at $ 2,675,000.

Book value per share = $ 2,675,000 / ( 59,000 x 1.16 ) = $ 2,675,000 / 68,440 shares = $ 39.085

Total book value for the stockholder holding 450 shares = 450 x 1.16 x $ 39.085 = $ 20,402.37